OpenStudy (anonymous):

What is the equation of the directrix of the parabola with equation x=4y^2+16y+19

OpenStudy (dumbcow):

first find the vertex by using completing the square

OpenStudy (anonymous):

ill just guess on this one thanks anyway

OpenStudy (dumbcow):

haha ok, don't you want to learn how to do these problems here i'll show you the steps: i'll use the shortcut way of getting vertex --> y = -b/2a = -16/8 = -2 plug that in to get x_coordinate x = 4(-2)^2 +16(-2) +19 = 3 vertex: (3,-2) directrix: x = h-p where p = 1/(4a) = 1/16 h = 3 --> x = 3- 1/16 --> x = 48/16 - 1/16 --> x = 47/16

OpenStudy (anonymous):

I just put x=-2 1/16

OpenStudy (dumbcow):

?

OpenStudy (anonymous):

for the answer. It is multiply choice so that is what i chose

OpenStudy (dumbcow):

yeah you guessed...but i showed you the solution, can you follow my steps?

OpenStudy (anonymous):

i cant learn over typing. I have to hear someone walk me through it. I have memory loss too so i wont be able to remember it anyway

OpenStudy (dumbcow):

oh ok so you probably should go somewhere else for help then anyway answer is x=47/16 or x=2 15/16

OpenStudy (anonymous):

i know i should but thank you for helping